OpenStudy (anonymous):

2x-7y=14 can someone break this down barney style for slope with the steps

OpenStudy (anonymous):

I know the formula I just don't understand exactly how to get it

OpenStudy (aum):

You have to put the equation in the form: y = mx + b. That means you have to solve the given equation for y. 2x-7y=14 To solve for y, isolate y: add 7y to both sides: 2x = 14 + 7y subtract 14 from both sides: 2x - 14 = 7y divide both sides by 7 2/7x - 2 = y or y = 2/7 * x - 2 Compare this to the general slope-intercept form y = mx + b. What is m?

OpenStudy (aum):

In y = mx + b, 'm' is the slope. Compare above equation to y = 2/7 * x - 2 we got earlier. What is m?

OpenStudy (anonymous):

is it undefined?

OpenStudy (aum):

y = mx + b y = 2/7x - 2 compare the two. In the first equation, m is the coefficient of x In the second equation, 2/7 is the coefficient of x. So m = 2/7. Therefore, the slope of the given equation is 2/7.

OpenStudy (anonymous):

okay but what about the -2